A 690 g of 59% strontium bromide solution is cooled from 60 oC to a certain temperature. Upon addition
of around 2 g of strontium bromide, excess solute above saturation started to crystallize. Total
recovered potassium bromide crystals is 146.003 g.

5a) How many solute is initially dissolved?
5b) In the recovered crystals, how many grams came from the solution?
5c) How many grams are still dissolved after crystallization?
1
Expert's answer
2020-10-22T14:19:10-0400

5a. (690×59)/100=407.1grams

5b. 146.003-2=144.003

5c. 407.1-144.003=263.097